OpenStudy (mikeymaximum):

Ms. Lunette has 18 students in her class. She wants to send 3 of these students to pick up books for the class. How many combinations of 3 students can she choose? A. 6 B. 54 C. 816 D. 4896

OpenStudy (mikeymaximum):

I know for sure that the answer is not either A or D

OpenStudy (mikeymaximum):

I think most likely C

OpenStudy (mikeymaximum):

tbh I'm not sure

imqwerty (imqwerty):

The number of ways of selecting 'r' number of different things from 'n' number of different things is nCr. So, here, Ms. Lunette has to select 3 different students from 18 different students. The Number of ways of doing so is 18C3 = 18!/3!(18-3)! = 18!/3!15! = (16 * 17 * 18) / (1 * 2 * 3) = 816. The answer is c.

OpenStudy (nincompoop):

can she repeat people?

OpenStudy (anonymous):

18C3 = 816

OpenStudy (mikeymaximum):

Yes She can

OpenStudy (nincompoop):

then follow qwerty's concept

OpenStudy (mikeymaximum):

ok Thx 2 all who came to help thou

imqwerty (imqwerty):

;)